# A Comprehensive Review on Harnessing Large Language Models to Overcome Recommender System Challenges

### v1

@@ 4. Evolving Recommender Systems: Addressing Traditional Challenges with LLM-Based Solutions
(Zhang et al., 2021a; Li et al., 2023a; Zhai, 2024; Xu et al., 2023b; Kang et al., 2023; Dai et al., 2023)
@@ 4.1.1. Cold Start Problem
This challenge is exacerbated in large-scale, dynamic environments where thousands of new users and items enter the system daily (Quadrana et al., 2017). Even more complex architectures such as two-tower models (Covington et al., 2016) or deep sequential recommenders (Hidasi et al., 2016) rely on embedding representations learned from prior interactions, and thus remain vulnerable to cold start scenarios. The sparsity of user–item signals and the lack of initialization paths for unseen entities make cold start a persistent bottleneck in the design of effective recommender systems (Schein et al., 2002; Park and Tuzhilin, 2006). LLMs offer a paradigm shift in mitigating the cold start problem by leveraging their ability to process rich textual and contextual information without requiring historical interaction data (Dai et al., 2023; Wang et al., 2024b). Unlike traditional collaborative filtering methods that depend on dense user–item matrices, LLMs can infer relevance through semantic reasoning, prompt-based generation, and cross-modal understanding. Let’s examine how LLMs can solve the cold start problem in recommender systems.
@@
Retrieval-Augmented Generation (RAG) – RAG-based methods enhance LLMs with a retrieval module that selects relevant items or documents based on the input query. In recommender systems, this architecture decouples memorization from inference, allowing the model to attend to similar items, behavioral patterns, or metadata even when the target entity is entirely new or sparsely observed (Kang et al., 2023; Dai et al., 2023). By combining parametric knowledge from the LLM with non-parametric access to a large external index, RAG frameworks enable more accurate, controllable, and interpretable recommendation outputs.
@@
Notably, recent systems like PromptRec (Dai et al., 2023), PROMO (Jiang et al., 2024), and OpenAI’s plugin-based RecSys prototype (Li et al., 2023a) have operationalized such strategies in both online and offline settings. These systems dynamically compose prompts using structured user/item attributes and contextual cues (e.g., time, location, browsing intent), effectively merging retrieval and generation within a single architecture.
@@
Multimodal Embedding Synthesis – When textual information alone is insufficient for high–fidelity personalization, LLMs can be situated at the center of a multimodal pipeline that fuses visual, structured, and behavioral signals into a unified representation  (Ranasinghe et al., 2025). A generic fusion template is:
@@
where the gating logits $g_{m}$ are produced by an attention-over-modalities layer conditioned on contextual cues (e.g. user intent or query). This adaptive weighting has two practical advantages: (i) it suppresses noisy modalities when signals are weak (e.g. blurry images) and (ii) it accentuates the most informative channel for each request, leading to more stable cold-start performance  (Yu et al., 2022). Amazon’s Multimodal RecSys (Tian et al., 2025) employs hierarchical fusion: fine-tuned CLIP encoders supply image vectors, BERT-based encoders supply text, and catalog metadata is featurized via embedding tables. A shallow cross-modal transformer merges these signals and outputs $\mathbf{q}_{i}$ to a traditional ANN service, yielding a $\sim$5–10lift in hit-rate for products with <5 historical interactions  (Xu et al., 2023c). Similar blueprints are emerging in streaming platforms that blend thumbnails, transcripts, and user watch-history to build rich item sketches for zero-shot retrieval  (Xu et al., 2022). Future work is exploring co-training regimes where the LLM simultaneously solves captioning  (Hao et al., 2024), visual question-answering, and recommendation; this multitask conditioning regularizes the shared space and mitigates over-fitting to any single modality. Knowledge-distillation and quantization pipelines are being introduced to compress the fused encoders for edge deployment, aligning multimodal personalization with latency and privacy constraints.
Meta-Learning for Fast Adaptation – Meta-learning provides a framework for enabling rapid personalization by training models to adapt quickly to new users or items using limited data  (Wang et al., 2024c). When combined with LLMs, meta-learning techniques such as Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) allow systems to fine-tune user- or item-specific representations via a small number of gradient updates. For a new user task $\mathcal{T}_{u}$, the model parameters are adapted as follows:
@@
Recent approaches such as PromptRec and LLMRec show that this strategy improves recommendation relevance in cross-modal and multilingual setups. Additionally, systems like CrossAligner (Gritta et al., 2022) use contrastive fine-tuning to align embeddings across domains, while  (Zhang et al., 2025c) demonstrate strong performance on cold-start item transfer using instruction-following LLMs. This line of work opens avenues for applying pretraining-era language representations to cross-domain personalization at scale, eliminating the need for costly alignment mappings or dual-domain retraining.
@@
(a) (i)
(b) (ii)
Figure 7. (i) LLM-Based Semantic Matching. The LLM converts textual metadata of new users and new items into dense embeddings. These embeddings are semantically matched via dot product to generate recommendation scores, enabling cold-start recommendations without historical interaction data. (ii) LLM-Based Personalization Techniques. The figure highlights four core strategies used by Large Language Models to enhance personalization: (1) Instruction-Tuned Generalization ensures global instruction-following; (2) Prompt-Tuned Personalization leverages embedded user feedback; (3) Behavioral Diversity is enabled through generative modeling; and (4) Multitask Prompting supports learning from multiple objectives. Together, these approaches enable more adaptive and user-aligned recommendation systems.
@@ 4.1.5. Multimodal Data Integration
Unified Representation through Instruction-Tuned LLMs – Given a structured prompt $\texttt{Prompt}_{i}$ that serializes all modality features (e.g., title, description, image caption, metadata, and recent user feedback), an instruction-tuned LLM can embed this multimodal input using a shared semantic space. The dense item representation is computed as:
@@
Cross-Modal Alignment via Pretraining – Large-scale pretraining on multimodal corpora (e.g., alt-text–image pairs, metadata–review co-occurrences) enables LLMs to implicitly align semantically related content across modalities  (Sapkota et al., 2025). These pretrained models induce a joint latent space where concepts are consistently grounded across different input forms  (Zhang et al., 2025d). At inference, when certain modalities are unavailable, the model can still generate meaningful representations by conditioning on the observed subset $x_{i}^{\text{obs}}\subset x_{i}$:
@@
This alignment strategy originated in vision-language models like CLIP  (Radford et al., 2021) and ALIGN  (Jia et al., 2021), and has been adapted for recommendation in PaLI (Chen et al., 2022), CoCa-2Rec (Wang et al., 2024e), and OmniFM (Zhou et al., 2025), where the model generalizes across image-text-review-metadata domains with minimal supervision.
@@ 4.2.2. Scalability of Deep Models
where $p_{\text{LLM}}$ denotes the teacher’s distribution and $p_{\text{student}}$ the distilled model. Companies like TikTok (Zhao et al., 2022), Amazon (Shen et al., 2021), and Alibaba (Li et al., 2023a) use LLM distillation to create small, latency-aware models that retain generative capabilities for recommendations, search, and ranking. While traditional DNNs can be distilled, LLM-based distillation uniquely preserves instruction-following, multi-domain reasoning, and natural language grounding. Moreover, LLMs support multi-task distillation (e.g., joint training on ranking, generation, and explanation), enabling compressed models to generalize across different RecSys objectives (Dai et al., 2023; Liu et al., 2023c).
@@
Sparse Activation Architectures – Sparse MoE (Mixture of Experts) models activate only a subset of parameters during inference, scaling to billions of weights without proportional cost (Lepikhin et al., 2021; Zoph et al., 2022). For recommender systems, LLM-based MoE allows selective activation based on user intent or domain — as shown in GLaM (Du et al., 2022), Switch-LLM (Fedus et al., 2022), and V-MoE (Riquelme et al., 2021), which uses sparsity to optimize vision transformer inference. Amazon’s M6-T (Lin et al., 2021) and DeepMind’s GopherMoE (Rae et al., 2021) extend this approach for multi-task retrieval, recommendation, and generation. This reduces effective inference complexity:
@@ 4.2.3. Long-Tail Modeling
where $\text{Rel}(i,u)$ denotes predicted user-item relevance, $\text{Pop}(i)$ is a popularity prior, and $\lambda\in[0,1]$ balances relevance against popularity bias. A small $\lambda$ skews the model towards head items, suppressing cold and long-tail entities.
@@
Retrieval-Augmented Tail Expansion – Tail items often suffer from cold-start issues due to minimal or no user engagement. Retrieval-Augmented Generation (RAG) frameworks alleviate this by fetching external content—such as similar item reviews, frequently co-viewed products, or question–answer forums—that offer surrogate interaction signals. These retrieved contexts act as priors, conditioning the LLM to produce recommendations or relevance scores with greater semantic grounding. Formally, the decoder incorporates $\mathcal{R}(x_{i})$, the retrieved context for item $x_{i}$, to better personalize outputs for user $x_{u}$. This paradigm has gained real-world traction in platforms like Amazon and Spotify (Wang et al., 2023b; Vasile et al., 2022), where hybrid systems interleave retrieval and generation in ranking. Public systems such as RAG (Lewis et al., 2021) and PromptRec (Dai et al., 2023) provide open-source implementations of this approach. RAG mechanisms also allow continual updates to item knowledge without re-training the LLM, offering scalability for dynamic catalogs.
@@ 4.3. Evaluation and Experimentation Challenges
Figure 10. Illustration of Offline-Online Metric Divergence. This figure highlights the gap between offline evaluation—using historical logged data and metrics like NDCG, Recall@K, and MAP—and online deployment performance with real users, which is measured through business-critical KPIs such as click-through rate (CTR), average revenue per user (ARPU), session duration, and long-term user retention. This divergence presents a key challenge in validating recommender systems effectively.
@@ 4.3.1. Offline-Online Gap
[…]
@@
Behavior-Level Satisfaction The user clicked $x\_1$, skipped $x\_2$, then bounced. How satisfying was this session?
@@ 4.3.2. Sparse Conversion Labels
Proxy Signal Augmentation via LLMs - When labeled conversions are scarce, LLMs can extrapolate soft labels from partial interaction data such as clicks, dwell time, or page visits. The generated score $\hat{y}_{ui}^{\text{soft}}=\sigma(\text{LLM}(x_{u},x_{i}))$ can serve as a training target in binary classification or ranking setups. A representative prompt is shown in Table 6. This is especially valuable in semi-supervised learning regimes, where high-coverage but noisy labels can bridge the sparsity gap and enhance downstream recall (Dai et al., 2023; Zhai, 2024; Zhang et al., 2021a).
@@ 4.3.3. Balancing Immediate Engagement with Long-Term User Value
[…]
@@
where each ^$\hat{\mathcal{R}}_{t}$ is predicted by the LLM under a hypothesized slate. User-simulator frameworks such as SimRec (Chen et al., 2023b) and UserGPT-Sim (Zheng et al., 2023) report strong correlations (0.82–0.87) between simulated and real A/B outcomes.
@@ 4.4.4. Data Access and Governance
Balancing innovation with data security requires robust data governance. Access control systems that prevent unauthorized access must coexist with tools that enable rapid development and experimentation. Companies like LinkedIn and Airbnb have built governance platforms that include role-based access control (RBAC), automated policy enforcement, and secure experimentation sandboxes (Liu et al., 2021). These systems help enforce compliance while supporting agility. Teams should adopt scalable governance frameworks that integrate access controls directly into ML and data infrastructure, ensuring accountability without obstructing productivity.
Data Minimization via LLMs: One of the most promising applications of LLMs is their ability to reduce dependence on personally identifiable information (PII) by extracting meaningful signals from anonymized or high-level interaction data. Instead of collecting demographic features or granular histories, LLMs can infer user interests through textual representations of behavior, public reviews, or short intent prompts. This reduces the risk surface area while maintaining personalization effectiveness.
The goal is to minimize the quantity of personal data needed for effective recommendation:
$Q(\text{Rec})\approx Q_{\text{LLM}}(\text{Rec}\mid\text{anonymized data})$
Here, $Q(\text{Rec})$ represents recommendation quality. If the quality from the LLM using anonymized inputs approximates that from full PII, the system has achieved successful data minimization.
Recent work has shown that privacy-preserving techniques can be successfully applied to large language models in recommendation settings. One promising direction involves federated prompt learning with differential privacy, which can achieve performance comparable to centralized training while ensuring strong user-level protections  (Tran et al., 2025). Another approach leverages synthetic user simulations generated by LLMs, reducing reliance on sensitive behavioral logs. Additional strategies include token-level obfuscation to filter private information before inference, and private fine-tuning methods that mitigate membership inference risks without significantly compromising model accuracy  (Zhang et al., 2025b).
Instructable Privacy Filters: LLMs can also be explicitly guided—via prompt templates or fine-tuning—to avoid generating sensitive or non-compliant outputs. For example, when asked to explain a recommendation, the model can be instructed not to include health-related inferences or user-specific demographics. These filters act as soft constraints that shape the language generation output toward safer and more compliant formats.
Such behavior can be formalized with a constraint function:
$P(\text{output contains PII})<\delta$
where $\delta$ is a defined privacy threshold (ideally close to 0). Achieving this objective requires robust prompt design, alignment with safety constraints, and comprehensive evaluation procedures  (Chen et al., 2025a). Recent advancements have explored methods such as instruction-driven privacy regularization, rule-guided decoding to suppress sensitive content, and generation pipelines explicitly constrained to avoid exposing personally identifiable information. These safeguards are especially critical for large-scale systems that produce user-facing outputs like explanations, summaries, or personalized recommendations in natural language  (Dong et al., 2024b)
Synthetic Data Generation for Privacy: LLMs can also generate entirely synthetic user behavior data—such as clickstreams, session logs, or user profiles—that mimic the statistical structure of real data without revealing any individual information. This synthetic data can be used for training, testing, or debugging recommender systems in compliance with privacy constraints.
To ensure this approach is effective, synthetic datasets $D_{\text{synth}}$ must satisfy differential privacy guarantees. A typical requirement is:
$\Pr[\text{LLM outputs }D_{\text{synth}}\mid x\in D]\approx\Pr[\text{LLM outputs }D_{\text{synth}}\mid x\notin D]$
This ensures that the inclusion or exclusion of any individual data point in the training corpus does not significantly influence the model’s outputs. Additional metrics like Jensen-Shannon divergence or Wasserstein distance can be used to compare the distributional similarity between real and synthetic data.
Recent work has explored differentially private synthetic sequence modeling using transformer-based LLMs (Xie et al., 2024), conditional user profile generation for safe recsys debugging, and adversarially trained privacy-preserving simulators for behavioral logs (Lin et al., 2024). These models enable privacy-safe benchmarking and model iteration without relying on production data.

### v2

@@ 4. Evolving Recommender Systems: Addressing Traditional Challenges with LLM-Based Solutions
(Zhang et al., 2021a; Li et al., 2023a; Zhai, 2024; Xu et al., 2023b; Kang et al., 2023; Dai et al., 2023b)
@@ 4.1.1. Cold Start Problem
This challenge is exacerbated in large-scale, dynamic environments where thousands of new users and items enter the system daily (Quadrana et al., 2017). Even more complex architectures such as two-tower models (Covington et al., 2016) or deep sequential recommenders (Hidasi et al., 2016) rely on embedding representations learned from prior interactions, and thus remain vulnerable to cold start scenarios. The sparsity of user–item signals and the lack of initialization paths for unseen entities make cold start a persistent bottleneck in the design of effective recommender systems (Schein et al., 2002; Park and Tuzhilin, 2006). LLMs offer a paradigm shift in mitigating the cold start problem by leveraging their ability to process rich textual and contextual information without requiring historical interaction data (Dai et al., 2023b; Wang et al., 2024b). Unlike traditional collaborative filtering methods that depend on dense user–item matrices, LLMs can infer relevance through semantic reasoning, prompt-based generation, and cross-modal understanding. Let’s examine how LLMs can solve the cold start problem in recommender systems.
@@
Retrieval-Augmented Generation (RAG) – RAG-based methods enhance LLMs with a retrieval module that selects relevant items or documents based on the input query. In recommender systems, this architecture decouples memorization from inference, allowing the model to attend to similar items, behavioral patterns, or metadata even when the target entity is entirely new or sparsely observed (Kang et al., 2023; Dai et al., 2023b). By combining parametric knowledge from the LLM with non-parametric access to a large external index, RAG frameworks enable more accurate, controllable, and interpretable recommendation outputs.
@@
Notably, recent systems like PromptRec (Dai et al., 2023b), PROMO (Jiang et al., 2024), and OpenAI’s plugin-based RecSys prototype (Li et al., 2023a) have operationalized such strategies in both online and offline settings. These systems dynamically compose prompts using structured user/item attributes and contextual cues (e.g., time, location, browsing intent), effectively merging retrieval and generation within a single architecture.
@@
Multimodal Embedding Synthesis –When textual information alone is insufficient for high–fidelity personalization, LLMs can be situated at the center of a multimodal pipeline that fuses visual, structured, and temporal signals  Figure 3). d behavioral signals into a unified representation  (Ranasinghe et al., 2025). A generic fusion template is:
@@
Figure 3. Overview of the Proposed Multi-Modal Sequence Learning Framework.(a) The model processes timestamp, ID, text, and image sequences using BERT, ViT, and ID embeddings, followed by Interactive and Temporal MoE modules for expert routing. Outputs are fed into Transformers for prediction. (b) Multi-task learning includes category prediction, contrastive learning on ID embeddings, and placeholder-based contrastive learning across modalities  (Zhang et al., 2025b).
@@
where the gating logits $g_{m}$ are produced by an attention-over-modalities layer conditioned on contextual cues (e.g. user intent or query). This adaptive weighting has two practical advantages: (i) it suppresses noisy modalities when signals are weak (e.g. blurry images) and (ii) it accentuates the most informative channel for each request, leading to more stable cold-start performance  (Yu et al., 2022). Amazon’s Multimodal RecSys (Tian et al., 2025) employs hierarchical fusion: fine-tuned CLIP encoders supply image vectors, BERT-based encoders supply text, and catalog metadata is featurized via embedding tables. A shallow cross-modal transformer merges these signals and outputs $\mathbf{q}_{i}$ to a traditional ANN service, yielding a $\sim$5–10% lift in hit-rate for products with <5 historical interactions  (Xu et al., 2023c). Similar blueprints are emerging in streaming platforms that blend thumbnails, transcripts, and user watch-history to build rich item sketches for zero-shot retrieval  (Xu et al., 2022). Future work is exploring co-training regimes where the LLM simultaneously solves captioning  (Hao et al., 2024), visual question-answering, and recommendation; this multitask conditioning regularizes the shared space and mitigates over-fitting to any single modality. Knowledge-distillation and quantization pipelines are being introduced to compress the fused encoders for edge deployment, aligning multimodal personalization with latency and privacy constraints.
Meta-Learning for Fast Adaptation – Meta-learning provides a framework for enabling rapid personalization by training models to adapt quickly to new users or items using limited data  (Wang et al., 2024c). When combined with LLMs, meta-learning techniques such as Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017a) allow systems to fine-tune user- or item-specific representations via a small number of gradient updates(see Figure 4. For a new user task $\mathcal{T}_{u}$, the model parameters are adapted as follows:
@@
Figure 4. The image compares MAML-based few-shot adaptation (left) to standard fine-tuning without MAML (right) for a sine wave regression task. MAML generalizes well, even in regions without data, by capturing the underlying periodic structure. In contrast, the pretrained model struggles to adapt from few examples, failing to extrapolate due to conflicting gradients from pretraining (Finn et al., 2017b)
@@
Recent approaches such as PromptRec and LLMRec show that this strategy improves recommendation relevance in cross-modal and multilingual setups. Additionally, systems like CrossAligner (Gritta et al., 2022) use contrastive fine-tuning to align embeddings across domains, while  (Zhang et al., 2025d) demonstrate strong performance on cold-start item transfer using instruction-following LLMs. This line of work opens avenues for applying pretraining-era language representations to cross-domain personalization at scale, eliminating the need for costly alignment mappings or dual-domain retraining.
@@ 4.1.5. Multimodal Data Integration
Figure 7. An instruction-tuned LLM framework that unifies visual and textual inputs into a shared semantic space. The image encoder extracts visual embeddings, which are fused with textual instructions via a Q-Former module using self-attention and cross-attention. The combined representation is processed by the LLM to generate a response, enabling cross-modal reasoning without separate modality-specific towers. (Dai et al., 2023a)
Unified Representation through Instruction-Tuned LLMs – Given a structured prompt $\texttt{Prompt}_{i}$ that serializes all modality features (e.g., title, description, image caption, metadata, and recent user feedback), an instruction-tuned LLM(see Figure 7 can embed this multimodal input using a shared semantic space. The dense item representation is computed as:
@@
Cross-Modal Alignment via Pretraining – Large-scale pretraining on multimodal corpora (e.g., alt-text–image pairs, metadata–review co-occurrences) enables LLMs to implicitly align semantically related content across modalities  (Sapkota et al., 2025). These pretrained models induce a joint latent space where concepts are consistently grounded across different input forms  (Zhang et al., 2025e). At inference, when certain modalities are unavailable, the model can still generate meaningful representations by conditioning on the observed subset $x_{i}^{\text{obs}}\subset x_{i}$:
@@
This alignment strategy originated in vision-language models like CLIP  (Radford et al., 2021) and ALIGN  (Jia et al., 2021), and has been adapted for recommendation in PaLI (Chen et al., 2022), CoCa-2Rec (Wang et al., 2024d), and OmniFM (Zhou et al., 2025), where the model generalizes across image-text-review-metadata domains with minimal supervision.
@@ 4.2.2. Scalability of Deep Models
where $p_{\text{LLM}}$ denotes the teacher’s distribution and $p_{\text{student}}$ the distilled model. Companies like TikTok (Zhao et al., 2022), Amazon (Shen et al., 2021), and Alibaba (Li et al., 2023a) use LLM distillation to create small, latency-aware models that retain generative capabilities for recommendations, search, and ranking. While traditional DNNs can be distilled, LLM-based distillation uniquely preserves instruction-following, multi-domain reasoning, and natural language grounding. Moreover, LLMs support multi-task distillation (e.g., joint training on ranking, generation, and explanation), enabling compressed models to generalize across different RecSys objectives (Dai et al., 2023b; Liu et al., 2023c).
@@
Figure 8. Illustration of a Sparse Mixture-of-Experts (MoE) layer where a gating network dynamically routes input tokens to a subset of experts. Only the top-k experts are activated per input, enabling efficient computation and scalability within Transformer architectures. (Shazeer et al., 2017)
Sparse Activation Architectures – Sparse Mixture of Experts(see Figure 8) models activate only a subset of parameters during inference, scaling to billions of weights without proportional cost (Lepikhin et al., 2021; Zoph et al., 2022). For recommender systems, LLM-based MoE allows selective activation based on user intent or domain — as shown in GLaM (Du et al., 2022), Switch-LLM (Fedus et al., 2022), and V-MoE (Riquelme et al., 2021), which uses sparsity to optimize vision transformer inference. Amazon’s M6-T (Lin et al., 2021) and DeepMind’s GopherMoE (Rae et al., 2021) extend this approach for multi-task retrieval, recommendation, and generation. This reduces effective inference complexity:
@@ 4.2.3. Long-Tail Modeling
where $\text{Rel}(i,u)$ denotes predicted user-item relevance, $\text{Pop}(i)$ is a popularity prior, and $\lambda\in[0,1]$ balances relevance against popularity bias. A small $\lambda$ skews the model towards head items, suppressing cold and long-tail entities. The following subsections discuss how long tail modeling can be solved using LLMs.
@@
Retrieval-Augmented Tail Expansion – Tail items often suffer from cold-start issues due to minimal or no user engagement. Retrieval-Augmented Generation (RAG) frameworks alleviate this by fetching external content such as similar item reviews, frequently co-viewed products, or question–answer forums that offer surrogate interaction signals. These retrieved contexts act as priors, conditioning the LLM to produce recommendations or relevance scores with greater semantic grounding. Formally, the decoder incorporates $\mathcal{R}(x_{i})$, the retrieved context for item $x_{i}$, to better personalize outputs for user $x_{u}$. This paradigm has gained real-world traction in platforms like Amazon and Spotify (Wang et al., 2023b; Vasile et al., 2022), where hybrid systems interleave retrieval and generation in ranking. Public systems such as RAG (Lewis et al., 2021) and PromptRec (Dai et al., 2023b) provide open-source implementations of this approach. RAG mechanisms also allow continual updates to item knowledge without re-training the LLM, offering scalability for dynamic catalogs.
@@ 4.3.1. Offline-Online Gap
$\tilde{r}_{ui}^{\text{LLM}}=\text{sigmoid}(\text{LLM}(x_{i},c_{u},\texttt{``How relevant is this item?''}))$
@@
Behavior-Level Satisfaction The user clicked x_1x\_1, skipped x_2x\_2, then bounced. How satisfying was this session?
@@ 4.3.2. Sparse Conversion Labels
Proxy Signal Augmentation via LLMs - When labeled conversions are scarce, LLMs can extrapolate soft labels from partial interaction data such as clicks, dwell time, or page visits. The generated score $\hat{y}_{ui}^{\text{soft}}=\sigma(\text{LLM}(x_{u},x_{i}))$ can serve as a training target in binary classification or ranking setups. A representative prompt is shown in Table 6. This is especially valuable in semi-supervised learning regimes, where high-coverage but noisy labels can bridge the sparsity gap and enhance downstream recall (Dai et al., 2023b; Zhai, 2024; Zhang et al., 2021a).
@@ 4.3.3. Balancing Immediate Engagement with Long-Term User Value
$\mathbf{u}_{\text{long}}=f_{\text{LLM}}(\texttt{user\_reply}),$
@@
where each $\hat{\mathcal{R}}_{t}$ is predicted by the LLM under a hypothesized slate. User-simulator frameworks such as SimRec (Chen et al., 2023b) and UserGPT-Sim (Zheng et al., 2023) report strong correlations (0.82–0.87) between simulated and real A/B outcomes.
@@ 4.4.4. Data Access and Governance
Balancing innovation with data security requires robust governance frameworks. Access control systems must prevent unauthorized use while enabling rapid experimentation. Companies such as LinkedIn and Airbnb employ platforms with role-based access control (RBAC), automated policy enforcement, and secure sandboxes (Liu et al., 2021). These ensure compliance and accountability without obstructing productivity.
Data Minimization via LLMs: LLMs can reduce reliance on personally identifiable information (PII) by extracting signals from anonymized data, reviews, or intent prompts, thereby minimizing privacy risks while maintaining personalization:
[…]
Privacy-preserving methods include federated prompt learning with differential privacy (Tran et al., 2025), synthetic user simulations, token-level obfuscation, and private fine-tuning (Zhang et al., 2025c).
Instructable Privacy Filters: LLMs can be guided through prompts or fine-tuning to avoid sensitive outputs, enforcing constraints such as:
[…]
where $\delta$ is a privacy threshold. Techniques include privacy regularization, rule-guided decoding, and instruction-driven alignment (Chen et al., 2025a; Dong et al., 2024b).
Synthetic Data Generation: LLMs can generate synthetic user logs or profiles that mimic real data distributions without exposing individuals. Differential privacy requires:
[…]
ensuring individual data points do not affect outputs. Metrics like Jensen-Shannon divergence or Wasserstein distance validate distributional similarity. Recent work explores DP-based sequence modeling (Xie et al., 2024), conditional profile generation, and adversarial simulators (Lin et al., 2024), enabling privacy-safe benchmarking and debugging.